# Evaluating LLMs’ Inherent Multi-hop Reasoning Ability

### v4

@@ 4.3 Ablation Study
To evaluate the impact of sub-questions for LLMs, we conduct an ablation study testing the performance of answering the final answer and removing the sub-questions from prompts. The results, shown in Table 4, indicate that when directly asking LLM a multi-hop question and corresponding passage, the performance is much lower than that of adding sub-questions to require LLMs reasoning step-by-step. For example, computed from table 4 the performance of GPT-4 on the 2-hop dataset decreased the F1 score and EM by 2.5 and 9.4 respectively. The results show the sub-questions could help LLMs improve the performance of final-QA.
4.4 Error Analysis
@@
We select a total of 20 incorrect final answers generated by GPT-4 from the 4-hop dataset to comprehensively illustrate how LLMs make decision on multi-step reasoning task. We first verify the proportion of each incorrect sub-answers and final-answers. Among the 20 incorrect final answers, 9 of them are wrongly answered in the first sub-questions that leads to the incorrect final answers. While for the remaining 11, 5 of them are incorrect second sub-questions that lead to the wrong final answers. The rest of 4 are influenced by the wrong third answers and fourth answers. From this analysis, we estimate that roughly half of the incorrect final answers are incorrectly reasoning from scratch. We also select 20 correct final answers generated by GPT-4 and find that about 4 of them are reasoned from incorrect reasoning chains (wrong sub-answers), revealing that LLMs also sometimes bypass the incorrect reasoning chain and get correct final answers.
4.5 Insights of LLM Multi-step Reasoning Evaluation
Drawing from LLMs’ multi-step reasoning performance and limitations and error analysis, we offer these insights:
Exact Matching
LLMs’ multi-step reasoning
Although the experiment results demonstrates that LLMs could perform multi-step reasoning ability in a certain extent, it remain sensitive to prompts and the impact of additional contexts, especially the sub-questions. Provide sub-questions as additional information into prompts can help guide LLMs to reason in a correct direction and show a relatively strong performance.
Reasoning chain Evaluation (Joint F1 RC and Joint EM RC in this work)
@@ 5 Conclusion
In this work, we present a novel evaluation method IRE which assesses the multi-step reasoning ability of LLMs via multi-hop QA and sub-QA and jointly computes the scores of the whole reasoning chain. To disentangle LLMs’ memory and reasoning ability, we design a framework that automatically synthesizes counterfactual data with human review to serve the evaluation. Although LLMs performed relatively well on QA tasks, the performance dropped on multi-hop questions that were based on new, counterfactual knowledge. In addition, their high performances are inflated and benefit from the high proportion of incorrect reasoning chains. We hope our work can facilitate future research on developing faithful knowledge editing methods.
@@ Appendix A Prompts and Examples
When evaluating large language models, prompting is a brittle process wherein small modifications to the prompt can cause large variations in the model predictions, and therefore significant effort should be dedicated to designing a painstakingly crafted perfect prompt for the given task [47, 48]. In this study, We investigate the performance of zero-shot on our benchmark. To eliminate the randomness, we manually select one demonstration for each task, ensuring that all tasks are covered.
We give our designed input examples for three different tasks to help readers understand our implementation, as shown in Table 7, respectively. The original and rewritten passages are shown in the table 9. The annotated multi-hop questions are shown in the table 8.
@@
Appendix B Joint Compuation
@@ Appendix C Performance Analysis
Table 10: The LLM evaluation on IRE 3 hop dataset. We here measure the sub-qa task and compare the performance between each hop. $Q_{i}$ means the $ith$ sub-questions.
@@
Table 11: The LLM performance on IRE 4 hop dataset. We here measure the sub-qa task and compare the performance between each hop.
@@
Figure 5: Performance gap between Wikipedia-based factual multi-hop QA datasets and our 2-hop, 3-hop, and 4-hop counterfactual MHQA data of table 3. The line charts reveal that LLMs show an obvious performance gap between previous datasets and our counterfactual QA data.
@@
Appendix D Limiatations

### v5

@@ 4.3 Impact of Sub-Question
To evaluate the impact of sub-questions for LLMs, we conduct an ablation study testing the performance of answering the final answer and removing the sub-questions from prompts. All LLMs are combined with Chain-of-Thought (Wei et al., 2022) settings, in which LLMs incrementally derive a series of intermediate steps. The results, shown in Table 6, indicate that when directly asking LLM a multi-hop question and corresponding passage, the performance is much lower than that of adding sub-questions to require LLMs reasoning step-by-step. For example, computed from table 6 the performance of GPT-4 on the 2-hop dataset decreased the F1 score and EM by 2.5 and 9.4 respectively. The results show the sub-questions could help LLMs improve the performance of final-QA.
5 Error Analysis
@@
We select a total of 20 incorrect final answers generated by GPT-4 from the 4-hop dataset to comprehensively illustrate how LLMs make decision on multi-step reasoning tasks. We first verify the proportion of each incorrect sub-answer and final answer. Among the 20 incorrect final answers, 9 of them are wrongly answered in the first sub-questions which leads to the incorrect final answers. While for the remaining 11, 5 of them are incorrect second sub-questions that lead to the wrong final answers. The rest of 4 are influenced by the wrong third answers and fourth answers. From this analysis, we estimate that roughly half of the incorrect final answers are incorrectly reasoning from scratch. We select 20 correct final answers generated by GPT-4 and find that about 4 of them are reasoned from incorrect reasoning chains (wrong sub-answers), revealing that LLMs also sometimes bypass the incorrect reasoning chain and get correct final answers. The insights of the whole experiment results are illustrated in Appendix D.
@@ 6 Conclusion
We presented a novel evaluation framework CoFCA for evaluating LLMs’ evidence integration and multi-step reasoning capabilities by combining sub-question evaluation and counterfactual data annotation. To disentangle LLMs’ memory and reasoning ability, we design a human-in-the-loop framework to synthesize counterfactual data. Our results show that, although LLMs performed relatively well on QA tasks, the performance dropped on multi-hop questions that were based on new, counterfactual knowledge. In addition, their high performances are inflated and benefit from a high proportion of incorrect reasoning chains. Our work can facilitate future research on developing faithful knowledge editing methods.
7 Reproducibility Statement
To make the results and models reproducible and verifiable, we provide our full data annotation guideline, data link, implementation details, and prompts: We detail the process of data annotation in section 3.1 and the implementations are in Appendix E. All the prompts required to reproduce the results are illustrated in Appendix A.
@@ Appendix A Prompts and Examples
When evaluating large language models, prompting is a brittle process wherein small modifications to the prompt can cause large variations in the model predictions, and therefore significant effort should be dedicated to designing a painstakingly crafted perfect prompt for the given task (Arora et al., 2022; Diao et al., 2023). In this study, We investigate the performance of zero-shot on our benchmark. To eliminate the randomness, we manually select one demonstration for each task, ensuring that all tasks are covered.
We give our designed input examples for three different tasks to help readers understand our implementation, as shown in Table 9, respectively. The original and rewritten passages are shown in the table 12.
The annotated multi-hop questions are shown in the table 11.
@@
Prompts of Partial Match Evalutaion
Prompt You are an Answer evaluator, you need to measure the semantic similarity between [Generated Answer] and [Gold Answer], and give the score, 1 means equal, 0 means not. Some answers may have abbreviations or alias, for example, Lionel Messi is equal to Messi, Donald Trump is equal to Trump. Please only output the score 1 or 0 without any other words. [Generated Answer] The LLM generated answer. [Gold Answer] The ground truth.
@@
Table 10: Examples of a passage annotated in different stages. The words in red indicate the revision part. Passage in stage 1 still has the same sentence structure with the original passage while passage in stage 2 is totally different.
Stage Passages HotpotQA Elliott Lester is an English film and television director, best known for directing the film "Blitz". He made his directing debut in 2006 with "Love Is the Drug", and his latest film, "Aftermath", was released on April 4, 2017. Stage 1 Marvin Kellaway is a British movie and radio helmsman, foremost noted for helming the movie ’Flashstrike’. He commenced his helming genesis in 2007 with ’Affection Is the Potion’, and his most recent movie, ’Repercussion’, was unveiled on May 3, 2018. Stage 2 Marvin Kellaway, a prominent British director in both film and radio, is best known for directing the popular movie ’Flashstrike’. He began his directing career in 2007 with ’Affection Is the Potion’, and his latest film, ’Repercussion’, was released on May 3, 2018.
@@
Appendix B Joint Computation
@@ Appendix C Performance Analysis
Table 13: The LLM evaluation on CofCA 3 hop dataset. We here measure the sub-qa task and compare the performance between each hop. $Q_{i}$ means the $ith$ sub-questions.
@@
Table 14: The LLM performance on CofCA 4 hop dataset. We here measure the sub-qa task and compare the performance between each hop.
@@
Figure 5: Performance gap between Wikipedia-based factual multi-hop QA datasets and our 2-hop, 3-hop, and 4-hop counterfactual MHQA data of table 3 and table 4. The line charts reveal that LLMs show an obvious performance gap between previous datasets and CofCA.
@@
Appendix D Insights of LLM Multi-step Reasoning Evaluation
Drawing from the above experimental results, we draw the conclusions:
Exact Matching
Multi-step Reasoning
Although the experiment results demonstrate that LLMs can perform multi-step reasoning ability to a certain extent, they remain sensitive to prompts and the impact of additional contexts, especially the sub-questions. Providing sub-questions as additional information into prompts can help guide LLMs to reason in the correct direction and show a relatively strong performance.
Reasoning chain Evaluation (Joint F1 RC and Joint EM RC in this work)
Appendix E Implementation Details
For proprietary models, we employ official APIs to interact with exclusive LLMs and prompts are well-defined. For open-source models, all experiments are conducted on 8 A100 GPUs.
Appendix F Limiatations